OpenStudy (anonymous):

The number of tickets sold each day for an upcoming performance of Handel's Messiah is given by N(x)=-0.4x^2+10.4x+13, where x is the number of days since the concert was first announced. When will daily ticket sales peak? and how many tickets will be sold that day?

OpenStudy (anonymous):

is it \[.4x^2+10.4x+13\]?

OpenStudy (anonymous):

find the vertex of the equation

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

if so max is given when \[x=\frac{-b}{2a}\] what mr orchestra said

OpenStudy (anonymous):

i get -13 but i must have read it wrong . oh wait

OpenStudy (anonymous):

\[-.4x^2+10.4x+13\]

OpenStudy (anonymous):

\[\frac{-b}{2a}=\frac{-10.4}{2\times -.4}=13\]

OpenStudy (anonymous):

then replace x by 13 and see what you get.

OpenStudy (anonymous):

24.62?

OpenStudy (anonymous):

11.61

OpenStudy (anonymous):

not what i got

OpenStudy (anonymous):

oops i forgot the - sign again!

OpenStudy (anonymous):

i got 80.6 let me try again

OpenStudy (anonymous):

got it twice

OpenStudy (anonymous):

I will keep trying I'm not sure what I'm doing wrong